Work Place Products Inc., a wholesaler of office products, was organized on July 1 of the current year, with an authorization of 100,000 shares of 3% preferred stock, $65 par and 450,000 shares of $15 par common stock. The following selected transactions were completed during the first year of operations:

Journalize the transactions.

July. 1. Issued 100,000 shares of common stock at par for cash.

July. 1

July. 1 Issued 650 shares of common stock at par to an attorney in payment of legal fees for organizing the corporation.

July. 1

Aug. 7. Issued 22,500 shares of common stock in exchange for land, buildings, and equipment with fair market prices of $61,000, $335,000, and $76,500, respectively.

For a compound transaction, if an amount box does not require an entry, leave it blank.

Aug. 7

Sept. 20. Issued 23,000 shares of preferred stock at $80 for cash.

For a compound transaction, if an amount box does not require an entry, leave it blank.

Sept. 20
Answer #1
July 1 Cash 1500000 =100000*15
      Common Stock 1500000
July 1 Organization expenses 9750 =650*15
      Common Stock 9750
Mar 9 Land 61000
Buildings 335000
Equipment 76500
      Common Stock 337500 =22500*15
      Paid in Capital in excess of par-Common Stock 135000
Sept 20 Cash 1840000 =23000*80
      Preferred Stock 1495000 =23000*65
      Paid in Capital in excess of par-Preferred Stock 345000